OpenStudy (anonymous):

what's (√3a^2b^2)(√9a^4b^6)(√12a^2)

pooja195 (pooja195):

@TheSmartOne

TheSmartOne (thesmartone):

Eh... Let me write the whole question properly so I can actually see what the actual question is :P

TheSmartOne (thesmartone):

\[(\sqrt{3a^2b^2})(\sqrt{9a^4b^6})(\sqrt{12a^2})\]

TheSmartOne (thesmartone):

Is that your question? :P

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

Well lets first start out with \(\bf\huge~~~~~~\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\bf ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~Made~by~TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

Now we can solve your question :P

OpenStudy (anonymous):

okay

TheSmartOne (thesmartone):

Lets first start off by simplifying the square roots :D What is this simplified? \((\sqrt{3a^2b^2})\)

OpenStudy (anonymous):

honestly, I really have no idea

TheSmartOne (thesmartone):

Well remember that anything that has \[\Large\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\]

TheSmartOne (thesmartone):

So lets do that with \((\sqrt{3a^2b^2})\)

TheSmartOne (thesmartone):

@ashleytomlinson99

OpenStudy (anonymous):

I feel really dumb for still not knowing how to do this

TheSmartOne (thesmartone):

No need to feel dumb. Lets do it step by step :) So we have \(\Large(\sqrt{3a^2b^2})\) and using \(\Large\sqrt{x}=x ^{\frac{ 1 }{ 2 }}\) we can try to remove the square root sign :D

TheSmartOne (thesmartone):

\(\Large (\sqrt{3a^2b^2})= (3a^2b^2)^\frac{1}{2}\)

TheSmartOne (thesmartone):

Do you get that part?

OpenStudy (anonymous):

yes

TheSmartOne (thesmartone):

So now lets use this formula :) \(\Large (abc)^\frac{1}{x}= a^\frac{1}{x}b^\frac{1}{x}c^\frac{1}{x}\)

OpenStudy (anonymous):

what I got was 1.7ab

TheSmartOne (thesmartone):

So what will \(\Large (3a^2b^2)^\frac{1}{2}\) be equal to?

TheSmartOne (thesmartone):

Correct but we can just to make life easier we can keep it as \(\large 3^\frac{1}{2}\)

OpenStudy (anonymous):

okay

TheSmartOne (thesmartone):

And we do the same process for the other two ones which are \(\Large (\sqrt{9a^4b^6})(\sqrt{12a^2})\)

TheSmartOne (thesmartone):

And also note that \(\Large 3^\frac{1}{2}= \sqrt{3}\)

OpenStudy (anonymous):

so then it's \[\sqrt{3ab}\] right

TheSmartOne (thesmartone):

Umm no. What is\(\Large (a^4b^6)^\frac{1}{2}\) The 3 part is correct :D

OpenStudy (anonymous):

okay it's \[a^{2} b ^{3}\]

TheSmartOne (thesmartone):

Yes! So now you got that \(\Large (\sqrt{9a^4b^6})= 3a^2b^3\)

OpenStudy (anonymous):

so I guess the 12 one would just be \[\sqrt{12}\] with a

TheSmartOne (thesmartone):

We could simplify 12 a little more down :D

TheSmartOne (thesmartone):

So what numbers when you multiply them equal to 12?

OpenStudy (anonymous):

\[\sqrt{4} \sqrt{3} = 2\sqrt{3}\]

TheSmartOne (thesmartone):

Correct! :D

TheSmartOne (thesmartone):

So now lets combine everything we have gotten so far!

OpenStudy (anonymous):

I know there is 4 a's and 4 b's

TheSmartOne (thesmartone):

\(\Large (\sqrt{3a^2b^2})(\sqrt{9a^4b^6})(\sqrt{12a^2})\) Was the original question. And we simplified it down to \(\Large \sqrt{3} ab \times 3 a^2 b^3 \times 2\sqrt{3}a\)

OpenStudy (anonymous):

so it's \[6a ^{4}b ^{4}\sqrt{3}\]

TheSmartOne (thesmartone):

\( a^4b^4\) is correct...

TheSmartOne (thesmartone):

What is \(\sqrt{3} \times \sqrt{3} =?\)

OpenStudy (anonymous):

3

TheSmartOne (thesmartone):

Correct

TheSmartOne (thesmartone):

and 3x3x2= ?

OpenStudy (anonymous):

\[18a ^{4}b ^{4}\]?

TheSmartOne (thesmartone):

Correct :D

TheSmartOne (thesmartone):

And that is your answer :)

OpenStudy (anonymous):

thank you that helps a lot. could you help me with a few more? I have 3

TheSmartOne (thesmartone):

Post a new question and tag @Nnesha and she can help you :) I have to go now :/

OpenStudy (anonymous):

alright thank you

TheSmartOne (thesmartone):

Or you can tag: mathmath333 sammixboo Luigi0210 jagr2713 sleepyjess And ask them to help you if nnesha doesn't come :)

TheSmartOne (thesmartone):

But Nnesha will always come :D